# GSCache: Real-Time Radiance Caching for Volume Path Tracing using 3D Gaussian Splatting

### v1

@@ 4.4 Runtime Performance
Runtime performance is an important factor for the use of a radiance cache as we generally want to achieve higher image quality in the same time or less compared to the uncached application.
@@ 5 Limitations and Future Directions
Cache Size. After initialization, during rendering and optimization, the cache resides entirely on the GPU. One downside of Gaussian-splatting-based approaches is the relatively high memory requirement compared to other scene representation techniques. This can negatively impact application performance as the cache might take up a non-negligible fraction of GPU memory during runtime. There have been several recent attempts to cull the memory footprint of Gaussian splatting applications [27, 21, 30, 35]. Future extensions of our approach could include such cache compression and reduction methods to create a smaller and more efficient cache. As a side-effect, such an optimization will likely result in some performance gain during runtime, as fewer Gaussians need to be optimized and rasterized.
Adaptivity Versus Consistency. For a radiance cache to be effective, it needs to react quickly to changes in the scene by evicting or forgetting past states and adapting to new ones. At the same time, if conditions are static, we want the cache to provide consistent data to ensure image quality and temporal stability between frames. In this work, we develop an adaptive learning rate scheduler that is tied to the renderer and user interaction patterns to determine the optimal learning rate at any given point. For our purposes, this constitutes an adequate trade-off between the adaptability and consistency of our cache. However, this method is specifically tailored to our scientific visualization use case and might not be adequate for other applications. In such cases, there is a need to develop more sophisticated methods to establish a balance between adaptability and consistency in the cache. Furthermore, while our design is largely agnostic to transfer function changes, it requires re-initialization for cases that alter the overall morphology of the visible volume. Future work could address refitting strategies for such cases. Another consequence of the cache’s high adaptivity its short-term memory. If a viewport is not covered for some time, the Gaussians encoding the radiance for this viewport might be re-used to encode more recent views, thus “forgetting” views that have not been active for some time and preventing a global convergence of the cache in the long term. While we did not observe unrecoverable cases of view-dependent overfitting, this limitation requires continuous training while scene or viewport changes are expected. Future work could address this issue by focusing optimizations on Gaussians that are closest and most relevant to the current viewport, potentially using spatial subdivision and nearest-neighbor methods to find optimization candidates to facilitate longer-term convergence of the cache.
@@
Self-Training. Currently, our cache is trained purely on unbiased path samples obtained during rendering. This limits training data to fully concluded paths. Future work could explore the potential of reusing cached entries for paths $\hat{n}<n$ by demodulating them with the throughput $\sigma_{i}$ and sampling probability $Tr$ of prior segments to generate samples for earlier cache levels in the spirit of Müller et al. [29].

### v2

@@ 4.4 Runtime Performance
Runtime performance is an important factor for the use of a radiance cache, as we generally want to achieve higher image quality in the same time or less compared to the uncached application.
@@ 5 Limitations and Future Directions
Cache Size. During rendering and optimization, the cache resides entirely on the GPU. One downside of Gaussian-splatting-based approaches is the relatively high memory requirement compared to other scene representation techniques. This can negatively impact application performance as the caching pipeline might take up a non-negligible fraction of GPU memory during runtime. There have been several recent attempts to cull the memory footprint of Gaussian splatting applications [31, 25, 35, 40]. Extensions of our approach could include such cache compression and reduction methods to create a smaller and more efficient cache.
Adaptivity Versus Consistency. For a radiance cache to be effective, it needs to react quickly to changes in the scene by evicting or forgetting past states and adapting to new ones. At the same time, if conditions are static, we want the cache to provide consistent data to ensure image quality and temporal stability between frames. In this work, we develop an adaptive learning rate scheduler that is tied to the renderer and user interaction patterns to determine the optimal learning rate at any given point. For our purposes, this constitutes an adequate trade-off between the adaptability and consistency of our cache. However, this method is specifically tailored to our scientific visualization use case and might not be adequate for other applications. In such cases, there is a need to develop more sophisticated methods to establish a balance between adaptability and consistency in the cache. Furthermore, while our design is largely agnostic to transfer function changes, it requires re-initialization for cases that alter the overall morphology of the visible volume. Future work could address refitting strategies for such cases. Another consequence of the cache’s high adaptivity is its short-term memory. If a viewport is not covered for some time, the Gaussians encoding the radiance for this viewport might be re-used to encode more recent views, thus “forgetting” views that have not been active for some time and preventing a global convergence of the cache in the long term. While we did not observe unrecoverable cases of view-dependent overfitting, this limitation requires continuous training while scene or viewport changes are expected. Future work could address this issue by focusing optimizations on Gaussians that are closest and most relevant to the current viewport, potentially using spatial subdivision and nearest-neighbor methods to find optimization candidates to facilitate longer-term convergence of the cache.
@@
Self-Training and Cache Design. Currently, our cache is trained in path space purely on unbiased path samples obtained during rendering. This limits training data to fully concluded paths. Future work could explore the potential of reusing cached entries for paths $\hat{n}<n$ by demodulating them with the throughput $\sigma_{i}$ and sampling probability $Tr$ of prior segments to generate samples for earlier cache levels in the spirit of Müller et al. [34]. Furthermore, recent advances in ray tracing Gaussian representations [5, 33, 54] allow for arbitrary ray queries which can enable world-space cache designs similar to NRC [34].
@@ Acknowledgements.
Appendix A Ablation Study
We provide an ablation study to showcase the effect of various choices regarding the initialization and training of our cache. In the following, we investigate how initialization size, spacing, and scaling of Gaussians, and various training hyperparameters influence the final quality of the images.
A.1 Cache Size
Figure 10: Comparison of the visual quality of our method when using different cache initialization sizes $N$ as shown on the FullBody. The columns show images captured from different runs at 102410241024 SPP using an initial cache size noted at the top of each column. Results show significant gains in quality up to around $N=300$k, after which additional gains become negligible.
In this study, we compare the image quality of our results when using different cache initialization sizes $N$. The results in Figure 10 indicate that from $N=10$k to about $N=300$k we see significant improvements in image quality. Beyond this size, the additional gains are marginal considering the increased processing and storage cost of higher numbers of Gaussians. This informed our choice of $N=300$k in our other experiments. Note that $N$ denotes the number of Gaussians in the first cache level and that subsequent levels follow the schedule we describe in the main document, in which each level is half the size of its predecessor. The lower cache sizes generally produce more washed-out looks in the final image. This is because fewer Gaussians are forced to represent larger areas of the image, which reduces local quality and adaptability to finer details. At $N=3$M, this effect becomes almost unnoticeable. We note that the exact cache size threshold needed to produce acceptable results can vary from dataset to dataset simply due to their size. The FullBody dataset shown in Figure 10 was one of the larger datasets in our tests, which is why we chose it for this ablation.
A.2 Hyperparameters
Figure 11: Ablation of our hyperparameter choice. The image compares results from our configuration with various modified versions of those settings. Data was captured by first rotating the camera around the object for 256 frames, followed by a recuperation period of another 256 frames in a still viewport. We present different combinations of settings with different choices removed from our final settings. The labels in the figure denote the following modifications. (-SC) removes the initial size cap on Gaussians. (-CO) removes our limitation on the scaling learning rate and allows scaling at a LR of 0.01250.01250.0125. Lastly, (-REG) removes the regularization during training.
To understand how our hyperparameter choice influences final image quality, we conduct an ablation over several key components in our pipeline. Figure 11 summarizes our findings. In this study, we test different combinations of hyperparameter settings, gradually removing constraints or replacing key components with alternatives. Specifically, we consider the removal of the size gap and downscaling of the initial set of Gaussians that we describe in the main document (Figure 11 (SC)). Furthermore, we allow optimization for Gaussian covariance (CO), which by default we disabled for training stability. Lastly, we test how the regularization influenced training stability by replacing AdamW [30] with a regular Adam [20] optimizer.
The results indicate that regularization has the highest impact on final cache quality. Without it, the cache fails to adapt to new viewports and gradients quickly explode, leaving us with numerous artifacts that make the cache practically unusable (see Figure 11 Ours vs. -REG). This is in contrast to Kerbl et al.’s [18] original implementation, which worked well without regularization. Allowing the optimization to scale Gaussians by setting the scaling LR to a non-zero value can help offset training instability in cases where the optimizer fails to do so (see Figure 11 -REG vs. -CO,-REG). We found this choice to make little difference when using regularization, and it is up to the specific application to determine the best setting. Finally, the initial size cap and downscaling of Gaussians helped preserve some of the finer details in the structure of the dataset while trading this advantage for a slightly noisier look (see Figure 11 Ours vs. -SC). Overall, the image quality with size cap and downscaling proved to be slightly better. In summary, we see that our choices worked well for our set of test data, but we acknowledge that their impact on the final image quality and training performance might differ depending on the specific application, and we encourage users to experiment with different settings. Aside from that, a valuable insight we derive is that without regularization, real-time optimization on noisy data does not seem to be a viable option.
Appendix B Memory Footprint
We briefly mention the memory footprint of our method in the main document. Here, we provide a more detailed analysis in Table 3. The cache components refer to the different parameters stored for Gaussian splats. All data is stored as 32-bit floating-point values, which leaves further room for optimization. All sizes are reported in megabytes.
Table 3: Breakdown of the memory footprint of our cache per component. Here $N$ is the size of each cache level, which is 300300300K, 150150150K, and  [^75] K, respectively, for our setup.
Comp. Dim. L0 Size L1 Size L2 Size
Position $N\times 3$ 3.433.433.43 MB 1.721.721.72 MB 0.860.860.86 MB
Rotation $N\times 4$ 4.584.584.58 MB 2.292.292.29 MB 1.141.141.14 MB
Color $N\times 3$ 3.433.433.43 MB 1.721.721.72 MB 0.860.860.86 MB
Scale $N\times 3$ 3.433.433.43 MB 1.721.721.72 MB 0.860.860.86 MB
Opacity $N\times 1$ 1.141.141.14 MB 0.570.570.57 MB 0.280.280.28 MB
Total $N\times 14$ 16.0216.0216.02 MB 8.018.018.01 MB 4.014.014.01 MB
Appendix C Cache Initialization Time
The cache is initialized at the beginning of the application by randomly sampling the volume to generate a point cloud that follows the density distribution of the volume and the chosen transfer function. To this end, we use the same sampling method [51] as in the main path tracer. In Table 4, we report average initialization timings per dataset. As our analysis shows, this process takes too much time to be repeated every frame, but it is not so slow as to prohibit re-initialization of the cache upon transfer function changes.
Table 4: Measurement of average initialization timings over five runs per dataset or a total of 303030 initializations with $N=300$K. The individual timing depends on the dataset size and average volume density as mapped by the transfer function.
Dataset Initialization Times [ms] Average Time [ms]
Carp 315.688315.688315.688, 328.575328.575328.575, 325.539325.539325.539, 309.039309.039309.039, 317.080317.080317.080 319.18319.18319.18
FullBody 415.402415.402415.402, 410.654410.654410.654, 416.495416.495416.495, 442.563442.563442.563, 453.676453.676453.676 427.76427.76427.76
MechanicalHand 296.847296.847296.847, 279.348279.348279.348, 305.291305.291305.291, 286.633286.633286.633, 290.239290.239290.239 291.67291.67291.67
Spider 314.366314.366314.366, 316.603316.603316.603, 310.582310.582310.582, 321.163321.163321.163, 307.771307.771307.771 314.10314.10314.10
Supernova 428.015428.015428.015, 440.562440.562440.562, 438.888438.888438.888, 432.853432.853432.853, 414.349414.349414.349 430.93430.93430.93
Zebrafish 684.607684.607684.607, 678.135678.135678.135, 675.923675.923675.923, 685.786685.786685.786, 691.880691.880691.880 683.27683.27683.27
Total - 411.15411.15411.15
Appendix D Image Quality
To complement the figures in the main document, we provide larger versions of the qualitative results in Figures 1213141516, and 17.
Figure 12: Visual quality of our method on the Carp dataset compared to the baseline path tracer. We show results for images at 1 SPP and compare our method (GSCache) against a baseline volume path tracer with uniform sampling (Uniform) and a version that uses next-event estimation (NEE).
Figure 13: Visual quality of our method on the FullBody dataset compared to the baseline path tracer. We show results for images at 1 SPP and compare our method (GSCache) against a baseline volume path tracer with uniform sampling (Uniform) and a version that uses next-event estimation (NEE).
Figure 14: Visual quality of our method on the MechanicalHand dataset compared to the baseline path tracer. We show results for images at 2 SPP and compare our method (GSCache) against a baseline volume path tracer with uniform sampling (Uniform) and a version that uses next-event estimation (NEE).
Figure 15: Visual quality of our method on the Spider dataset compared to the baseline path tracer. We show results for images at 1 SPP and compare our method (GSCache) against a baseline volume path tracer with uniform sampling (Uniform) and a version that uses next-event estimation (NEE).
Figure 16: Visual quality of our method on the Supernova dataset compared to the baseline path tracer. We show results for images at 1 SPP and compare our method (GSCache) against a baseline volume path tracer with uniform sampling (Uniform) and a version that uses next-event estimation (NEE).
Figure 17: Visual quality of our method on the ZebraFish dataset compared to the baseline path tracer. We show results for images at 1 SPP and compare our method (GSCache) against a baseline volume path tracer with uniform sampling (Uniform) and a version that uses next-event estimation (NEE).